Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 383     Entry time: 03 Oct 2023, 22:58
Author: Harald Sinn 
Group: XRO 
Subject: Shift end 

End of Euler shift: all 3 SASE2 mirrors are now working with the Euler MDL.

We left the beamline aligned to HED, with M1,M2 on Pt for upcoming HED experiment at 18 keV.

Pre-absorber closed, CRL1 out, Solid attenuater with some diamonds.

ELOG V3.1.4-7c3fd00